# LOGen: Toward LiDAR Object Generation by Point Diffusion

### v2

@@ Abstract
The generation of LiDAR scans is a growing area of research with diverse applications to autonomous driving. However, scan generation remains challenging, especially when compared to the rapid advancement of 2D and 3D object generation. We introduce a novel task: LiDAR object generation, requiring models to produce 3D objects as viewed by a LiDAR scan. This task focuses LiDAR scan generation on the most interesting aspect of scenes, the objects, while also benefiting from advancements in 3D object generative methods. We introduce a novel diffusion-based model to produce LiDAR point clouds of dataset objects, including intensity, and with an extensive control of the generation via conditioning information. Our experiments on nuScenes show the quality of our generations measured with new 3D metrics developed to suit LiDAR objects.
Figure 1: Novel LiDAR objects generated by LOGen using the conditioning information of real objects from the nuScenes validation set. Top row shows a bicycle (car on bottom row) generated via interpolations of conditioning information. Relative to the sensor, $rot$ is the object’s rotation while $d$ is the distance. Color is according to LiDAR intensity. More examples in Appendix.
@@ 1 Introduction
As 2D generative methods have exploded in their capacity to generate diverse and high quality images [32], one area of focus for 3D object generation has been understudied: objects in the context of LiDAR sensors. LiDAR is an essential modality for autonomous agents in driving scenarios. It provides accurate 3D information about the environment while being robust to adverse light conditions, making it crucial for safe navigation. The utility and corresponding shortcomings of LiDAR datasets [21, 28] makes the generation of LiDAR objects interesting from a variety of perspectives (sensor level generation for simulation [26], rare class augmentation [54], out-of-distribution generation [22], corner case generation [19], object disentanglement [30] etc.).
Objects perceived by LiDAR sensors are distinct from classically rendered 3D objects in the impact that the relative angle and distance of objects to the sensor has on the final appearance of an object [57]. While generated 3D meshes are uniform volumes, LiDAR objects have unique appearances depending on the intersection of the beam and the visible face of the object. Further, LiDAR objects contain an additional fourth channel of information when compared to classic 3D objects: the intensity of the laser beam reflected back to the sensor. Models which generate LiDAR objects must be capable of using this relative positioning and creating the intensity channel in order to generate realistic objects [46].
LiDAR object generation thus explores the latent capability of generative models to understand complex 3D relationships between objects, as capturing the resulting appearance of an object via a LiDAR sensor requires understanding the interplay of angle, distance, and orientation between the sensor and the object. Generating objects utilizing this geometric view conditioning is a more challenging evaluation of 3D generative methods [42].
@@
We introduce a transformer-based diffusion model, dubbed LOGen (for LiDAR Object Generator), which is designed to generate LiDAR scans of new objects. The generation is parameterized by the object class, angle, and distance from which objects are seen by the LiDAR sensor. This conditioning allows fine grained control of the appearance of the generated objects.
To the best of our knowledge there is no previous work on this topic. As we are the first to tackle this task, we compare LOGen against suitable alternatives that we design, inspired from other tasks: a sparse 3D CNN (MinkUNet [6]); a 2D transformer-based generator (PixArt-$\alpha$ [5]) adapted to 3D; and a 3D transformer-based generator (DiT-3D [27]), originally designed for traditional 3D objects, which we adapt to LiDAR objects. We show the quality of our generated objects compared to these baselines. We also evaluate at scene level on the semantic segmentation task to assess the realism of objects we generate.
Our contributions are as follows:
  • We are the first to propose a diffusion-based method specifically designed to produce LiDAR point clouds of objects, including intensity and with an extensive control of the generation (conditioning).
@@
  • Our experiments on nuScenes [3] show that despite the compact size of our model (just 7.5m parameters), we achieve superior quality, which we also measure with new metrics more suited to LiDAR objects.
We believe that the LiDAR object generation task is ripe for future exploration, and the promising results we present here will serve as a basis for this task.
@@ 2 Related Work
Various techniques have been proposed for 3D object generation: GANs [1, 37], VAEs [20, 44], normalizing flows [52], diffusion at point level [25, 42], and diffusion on features at voxel level with a transformer architecture [27]. However, these methods are only typically evaluated on datasets such as ShapeNet [4], ModelNet [48], or CO3D [35] after sampling points uniformly on shape surfaces, rather than following LiDAR sampling patterns. Their behavior with uneven point distributions of (sensor-specific) LiDAR scans is unexplored.
These methods do not use geometric conditioning related to viewpoint and do not generate point intensities. Further, most of these methods are trained and evaluated with a high and fixed number of points: typically 2048. It is not clear how they would behave with varying and sparse point clouds, as is the case of LiDAR objects. Only PointFlow intrinsically supports the generation of an arbitrary number of points, but it is also trained and evaluated on 2048 points [52].
LiDAR scan generation.
@@
Methods have been proposed to generate full LiDAR scans, by LiDAR simulation [11] in synthetic environments [53], or by learning to generate LiDAR data using energy-based models [57], VQ-VAEs [51], diffusion in range images [29, 15, 34] or diffusion on 3D points [31], possibly with images [55]. Further, simulation and learning can be combined to improve realism [26, 58]. As our work focuses on objects from existing datasets, scene-level generation serves as a source of inspiration but is not a direct comparison.
Some scan generation methods can be conditioned on a semantic map [34], or can place learned codes of object classes into a canvas representing the scene before decoding [51]. But this instance-level conditioning offers little control on object generation and it does not provide a way to augment an existing point cloud. [47] introduces text based conditioning in LiDAR scan generation, but does not evaluate the fidelity of the conditional generations, providing only qualitative evaluations of generated objects in scenes.
LiDAR object simulation.
LiDAR simulation on manmade CAD objects has been proposed to add variation to objects inserted into scenes [8, 9]. However these approaches are rigid in that CAD shapes have to match the object shape distribution in the region of the world where the original scans were captured. Moreover, this approach does not handle intensity: it only produces 3D point coordinates, not intensities. LiDARSim [26] tackles the problem of LiDAR simulation from a data driven approach similar to LOGen, however still more broadly at the scene level. They take a physics based approach to simulate LiDAR, then learn the resulting raydrop to produce realistic scans including intensity. Our approach directly learns object appearance and intensity from the data at an object level, eliminating the need for physics based simulation and collecting 3D objects.
LiDAR object generation.
To the best of our knowledge, only [49] considers the specific question of LiDAR object generation. The paper actually proposes five methods, based on a Gaussian mixture model (GMM), a variational auto-encoder (VAE), a latent generative adversarial network (L-GAN), an adversarial auto-encoder (AAE), and point cloud diffusion based on U-Net (PCD) for which no code is available. All these methods have significant limitations. First, they offer no conditioning, and generated points have no dimension nor intensities. Training objects must have 2048 points: objects with more points are downsampled, objects with less points get duplicated points, and objects with less than 256 points are ignored, while we go down to as few as 20 points. Evaluation is only at object level, with a classification task, not at scene level, and only 5 categories are evaluated, some with merged classes to ensure “sufficient granularity” (e.g., motorcycles with bicycles), while we do it on all 10 object classes of nuScenes.
@@ 3 Method
The goal of the LiDAR object generation task is to learn, for known classes in a labeled LiDAR dataset, to generate 3D point clouds of novel objects, as they would be captured by the same LiDAR sensor from arbitrary viewpoints. The training distribution is given by real-world objects extracted from 3D scans.
@@ 3.1 Object parameterization
[…]
In the end, the data distribution to learn consists of objects represented by a set of points $s=\{P_{i}\}_{i=1}^{N}$ in $\mathbb{R}^{N\times 4}$ with (centered) Cartesian coordinates and intensities. Moreover, we condition the generation on the following box information: $\kappa=(\phi,d,z,l,w,h)$. Class $c$ is not directly included as we train one model per class.
We do not normalize the point coordinates inside the object box, e.g., in [0,1]. This preserves the absolute dimensions of objects and the specificity of the LiDAR scanning patterns, depending in particular on the distance to the sensor. Consequently, the diffusion model is not constrained to generate points only inside the box. Experiments however show that only a few generated points sometimes fall out of the box, but not very far.
@@
To address this issue, we introduce a simple embeddings based on PointNet [33]. These features are just the output of a single 1D convolution, with which we project from our input dimension (3+1) into the embedding dimension. Because the point clouds (averaging from 40 points for bikes to 463 for bus) are in general magnitudes smaller than the point clouds generated in the original DiT-3D paper (thousands of points), we can use standard memory-efficient attention layers and avoid the need for voxelization. In the experiment section, we examine the impact of both types of embeddings.
@@
3.2 LiDAR object diffusion
Given a dataset with a number of object point clouds $S\,{=}\,\{s_{j}\}_{j=1}^{M}$ alongside conditions $K\,{=}\,\{\kappa_{j}\}_{j=1}^{M}$, the objective is to learn the conditional distribution $p(S|K)$. This is done via a Denoising Diffusion process: we learn to denoise input 4D Gaussian noise ($N$ noisy points) into a target shape ($N$ points representing a scanned object).
We rely on the time-discrete formulation of the diffusion process known as Denoising Diffusion Probabilistic Models (DDPM) [14], which is also used in DIT-3D [27] and in LiDAR scene generation methods [31]. We only sketch this formulation below. More details can be found in [14, 25] and in Appendix.
@@ 3.2 LiDAR object diffusion
The reverse process is provided by a denoising network with parameters $\theta$. The goal is to learn a distribution of the form $p_{\theta}(s_{t-1}|s_{t})=\mathcal{N}(s_{t-1}|\bm{\mu}_{\theta}(s_{t},t),\sigma^{2}_{t}\mathbf{I})$. This is learned via optimizing the evidence variational lower bound, that simplifies to minimizing the Kullback-Leibler divergence between the forward and reverse processes. After simplification, it boils down to minimizing a mean-squared loss between the model output $\hat{\epsilon}_{\theta}(s_{t},t)$ and the ground-truth Gaussian noise $\epsilon_{0}$. Given $\hat{\epsilon}_{\theta}(s_{t},t)$, we can then compute $\bm{\mu}_{\theta}(s_{t},t)$ and get a denoising step recurrence equation defining $s_{t-1}$ from $s_{t}$ (cf. Appendix).
A training step consists of sampling a different random time step $t$ for each element of the batch, producing the noised samples, and using the denoising neural network to predict the output of the reverse process.
To exploit conditioning, the above formulation is extended with the extra input of a condition $\kappa$, leading to a denoising network $\hat{\epsilon}_{\theta}(s_{t},t,\kappa)$. To apply the conditional guidance, we follow the Classifer Free Guidance formulation from [13]. This implies the dropout of conditioning information during the training procedure, where all conditions are replaced by a pre-defined null embedding. Using the null embedding results in an unconditioned pass of the model. A parameter $\lambda$ controls the strength of the guidance. The final predicted noise for a time step $t$ is computed as: ${\epsilon}_{\theta}(s_{t},t,\kappa)=\lambda\hat{\epsilon}_{\theta}(s_{t},t,\kappa)+(1-\lambda)\hat{\epsilon}_{\theta}(s_{t},t,\emptyset)$.
Figure 2: Architecture of LOGen. For architectures of our baselines PixArt-L and DiT-3DL please refer to Appendix.
@@ 3.3 Architecture
DiT-3D [27] is straightforwardly derived from DiT [32], and thus shares a similar architecture. PixArt-$\alpha$ [5] adds cross attention to DiT to inject conditions. Our baselines DiT-3DL and PixArt-L are directly adapted from their original counterparts to handle: (i) LiDAR data, i.e., point embedding, and (ii) LiDAR object conditioning. LOGen is then a modification of PixArt-L, as described below. The conditioning $\kappa=(\phi,d,z,l,w,h)$ is applied in the denoising network to direct the generative process. Given the importance of the conditioning in our task, we explore three variants in the architecture design space.
@@
  • AdaLN-single is the guidance proposed and used in PixArt-$\alpha$ [5]. This guidance relies on the addition of cross-attention layers which attend between the embedded noisy points and the conditioning information. While in the adaptive layer norm the conditions are always used to learn global shift and scale parameters, cross-attention layers allow the model to learn the guidance on a local scale. We do the same in PixArt-L.
  • We observe that the separation of the self-attention layers from the cross-attention layers by a scale operation in the PixArt-L architecture degrades the generation results (cf. Tab. 1 and 2). Thus we introduce a third type of conditioning in LOGen blocks, where we apply the cross attention immediately after the self-attention layer, and scale after these operations. The full LOGen architecture is shown in Fig. 2.

### v3

@@ Abstract
The generation of LiDAR scans is a growing topic with diverse applications to autonomous driving. However, scan generation remains challenging, especially when compared to the rapid advancement of image and 3D object generation. We consider the task of LiDAR object generation, requiring models to produce 3D objects as viewed by a LiDAR scan. It focuses LiDAR scan generation on a key aspect of scenes, the objects, while also benefiting from advancements in 3D object generative methods. We introduce a novel diffusion-based model to produce LiDAR point clouds of dataset objects, including intensity, and with an extensive control of the generation via conditioning information. Our experiments on nuScenes and KITTI-360 show the quality of our generations measured with new 3D metrics developed to suit LiDAR objects. The code is available at https://github.com/valeoai/LOGen.
Figure 1: Novel LiDAR objects generated by LOGen trained on the nuScenes train set, produced using the conditioning information of real objects from the nuScenes validation set. Relative to the sensor, rotrot is the object’s rotation while $d$ is the distance. Point color is according to LiDAR intensity. More examples are in the appendix.
@@ 1 Introduction
As 2D generative methods have exploded in their capacity to generate diverse and high quality images [35], one area of focus for 3D object generation has been understudied: objects in the context of LiDAR sensors. LiDAR is an essential modality for autonomous agents in driving scenarios. It provides accurate 3D information about the environment, making it crucial for safe navigation. The utility and corresponding shortcomings of LiDAR datasets [24, 31] makes the generation of LiDAR objects interesting from a variety of perspectives: sensor level generation for simulation [29], rare class augmentation [57], out-of-distribution generation [25], corner case generation [21], object disentanglement [33], etc.
Objects perceived by LiDAR sensors are distinct from classically rendered 3D objects in the impact that the relative angle and distance of objects to the sensor has on the final appearance of an object [60]. Further, LiDAR objects contain an additional channel of information when compared to classic 3D objects: the intensity of the laser beam reflected back to the sensor. Models which generate LiDAR objects must be capable of using this relative positioning and creating the intensity channel in order to generate realistic objects [49].
@@
We introduce a transformer-based diffusion model, dubbed LOGen (for LiDAR Object Generator), which is designed to generate LiDAR scans of new objects. The generation is parameterized by the object class, angle, and distance from which objects are seen by the LiDAR. This conditioning allows fine grained control of the appearance of generated objects.
We compare LOGen against suitable alternatives that we design, inspired from other tasks: a sparse 3D CNN (MinkUNet [8]); a 2D transformer-based generator (PixArt-$\alpha$ [7]) adapted to 3D; and a 3D transformer-based generator (DiT-3D [30]), originally designed for traditional 3D objects, which we adapt to LiDAR objects. We also compare to concurrent work [54]. We show the quality of our generated objects compared to these baselines. We also evaluate at scene level on the semantic segmentation task to assess the realism of objects we generate. Examples of generation are shown in Fig. 1. Our contributions are as follows:
  • We propose a diffusion-based method specifically designed to produce LiDAR point clouds of objects, including intensity and with an extensive control of generation (conditioning).
@@
  • Our experiments show that despite the compact size of our model (7.5M parameters), we achieve superior quality, which we also measure with new metrics suited to LiDAR objects.
We believe that the LiDAR object generation task is ripe for future exploration, and that the promising results we present here will serve as a basis for this task.
@@ 2 Related Work
3D object generation. Various techniques have been proposed for 3D object generation: GANs [1, 40], VAEs [22, 47], normalizing flows [55], diffusion at point level [28, 45], and diffusion on features at voxel level with a transformer architecture [30]. However, these methods are only typically evaluated on datasets such as ShapeNet [6], ModelNet [51], or CO3D [38] after sampling points uniformly on shape surfaces, rather than following LiDAR sampling patterns. Their behavior with uneven point distributions of (sensor-specific) LiDAR scans is unexplored. Besides, these methods do not use geometric conditioning related to viewpoint and do not generate point intensities. Further, most of these methods are trained and evaluated with a high and fixed number of points: typically 2048. It is not clear how they would behave with varying and sparse point clouds, as is the case of LiDAR objects.
@@
LiDAR data generation. Methods have been proposed to generate full LiDAR scans, by LiDAR simulation [13, 10, 11] in synthetic environments [56], or by learning to generate LiDAR data using energy-based models [60], VQ-VAEs [53], diffusion in range images [32, 17, 37, 4] or diffusion on 3D points [34], possibly with images [58]. Further, simulation and learning can be combined to improve realism [29, 61]. Some scan generation methods incorporate objects via: conditioning on a semantic map [37]; placing learned codes of object classes into a canvas representing the scene before decoding [53]; in-painting objects into range images [4]; and using text-based conditioning [50]. However these methods do not focus on the visual fidelity of generated objects, do not directly learn object appearance and intensity from the data at an object level, and do not evaluate the object perceptive quality.
OLiDM [54] is a concurrent work on LiDAR object generation that highlights the importance of distinguishing between foreground and background objects, employing a diffusion transformer model to generate LiDAR objects. They allow conditioning on object viewpoint parameters to produce localized objects, and perform scene level generation using the generated objects as conditions. LOGen in contrast focuses on architecture level decisions and uses novel methods of incorporating conditioning information and encoding points. LOGen features more robust evaluation of generated objects, with classwise metrics and perceptual evaluations including generated intensities.
@@ 3 Method
The goal of the LiDAR object generation task is to learn, for known classes in a labeled LiDAR dataset, to generate 3D point clouds of novel objects, as they would be captured by the same LiDAR sensor from arbitrary viewpoints. As opposed to simulation, the training distribution is given by real-world objects extracted from 3D scans.
@@ 3.1 Object parameterization
We also decrease the number of parameters regarding the box center location and orientation as the object appearance, seen from the sensor, only depends on the distance $d$ to the object and on the observation angle $\phi$, i.e., the angle between the object heading $\psi$ and the ray from the object bounding box center to the sensor. (As an illustration, $\phi\,{=}\,0$ means the object is facing the sensor, regardless of the object location around the ego vehicle.) In this conversion from Cartesian to cylindrical coordinates, the altitude $z$ of the box center is kept.
In the end, the data distribution to learn consists of objects represented by a set of points $s=\{P_{i}\}_{i=1}^{N}$ in $\mathbb{R}^{N\times 4}$ with (centered) Cartesian coordinates and intensities. We do not normalize the point coordinates to preserve the absolute dimensions of objects and the specificity of the LiDAR scanning patterns, depending in particular on the distance to the sensor.
Last, we condition the generation on the following box information: $\kappa=(\phi,d,z,l,w,h)$. While our generation is classwise, class $c$ is not directly included here as we train one model per class. This choice was made for compute and efficiency reasons: the per sample compute of single vs multi-class models is equivalent, thus allowing us to parallelize training. We nevertheless also experimented with a multi-class model, which was lagging behind the single-model-per-class approach in realism metrics (see the appendix).
@@
To address this issue, we introduce a simple embeddings based on PointNet [36]. These features are just the output of a single 1D convolution, with which we project from our input dimension (3+1) into the embedding dimension. Because the point clouds are in general magnitudes smaller (e.g., averaging from 40 points for bikes to 463 for bus in nuScenes) than the point clouds generated in the original DiT-3D paper (thousands of points), we can use standard memory-efficient attention layers and avoid the need for voxelization. In the experiment section, we examine the impact of both types of embeddings.
@@
Figure 2: Architecture of LOGen and baselines PixArt-L and DiT-3DL.
LiDAR object diffusion.
Given a dataset with a number of object point clouds $S\,{=}\,\{s_{j}\}_{j=1}^{M}$ alongside conditions $K\,{=}\,\{\kappa_{j}\}_{j=1}^{M}$, the objective is to learn the conditional distribution $p(S|K)$. This is done via a Denoising Diffusion process: we learn to denoise input 4D Gaussian noise ($N$ noisy points) into a target shape ($N$ points representing a scanned object, with intensities).
We rely on the time-discrete formulation of the diffusion process known as Denoising Diffusion Probabilistic Models (DDPM) [16], which is also used in DIT-3D [30] and in LiDAR scene generation methods [34]. During training, we sample a random time step $t$ and produce, via a forward noising process, a noisy sample $s_{t}$ of an input object point cloud $s_{0}=s$. Starting from a real LiDAR object $s_{0}$, the forward noising procedure produces a series of increasingly noisy 4D point clouds $s_{1},\ldots,s_{t}$, finally arriving at $s_{T}\,{\sim}\,\mathcal{N}(\mathbf{0},\mathbf{I})\,{\in}\,\mathbb{R}^{4}$. The training objective is to learn a distribution of the form $p_{\theta}(s_{t-1}|s_{t})=\mathcal{N}(s_{t-1}|\bm{\mu}_{\theta}(s_{t},t),\sigma^{2}_{t}\mathbf{I})$, parameterized by a denoising neural network.
@@
To exploit conditioning, the above formulation is extended with the extra input of a condition $\kappa$, leading to a denoising network $\hat{\epsilon}_{\theta}(s_{t},t,\kappa)$. To apply the conditional guidance, we follow the Classifer Free Guidance formulation from [15]. This implies the dropout of conditioning information during the training procedure, where all conditions are replaced by a pre-defined null embedding, which results in an unconditioned pass of the model.
@@ 3.2 Architecture
We now describe the architecture of LOGen and of our baselines, that we illustrate on Fig. 2. DiT-3D [30] is straightforwardly derived from DiT [35], and thus shares a similar architecture. PixArt-$\alpha$ [7] adds cross attention to DiT to inject conditions. Our baselines DiT-3DL and PixArt-L are directly adapted from their original counterparts to handle: (i) LiDAR data, i.e., point embedding, and (ii) LiDAR object conditioning. LOGen is then a further modification of PixArt-L, as described below. The conditioning $\kappa=(\phi,d,z,l,w,h)$ is applied in the denoising network to direct the generative process. Given the importance of the conditioning in our task, we explore three variants in the architecture design space.
@@
  • AdaLN-single is the guidance proposed and used in PixArt-$\alpha$ [7]. While in the adaptive layer norm the condition is used to learn global shift and scale parameters, cross-attention layers allow the model to learn the guidance on a local scale. We do the same in PixArt-L.
  • We observe that the separation of the self-attention layers from the cross-attention layers by a scale operation in the PixArt-L architecture degrades the generation results (cf. Tab. 1). Thus we introduce a third form of conditioning in LOGen blocks, where we apply the cross attention immediately after the self-attention layer, and scale after these operations.
@@ 4 Experiments
4.1 Experimental Setup
Datasets.
We evaluate our approach on nuScenes [5], which was acquired with a 32-beam LiDAR, thus creating notably sparse sweep patterns. To construct object samples to learn from, we use the semantic point labels and the 3D box annotations of objects. We take as object instance all points falling into a box, filtering out possible points of a different class. We also discard objects with less than 20 points. Point intensity, which ranges in 0–255, has a highly irregular distribution, with most examples clustered from 0–30, and with outliers. To create a more uniform intensity distribution, we use log max scaling.
In total, we extract 268,684 objects with at least 20 points, which covers more than 95% of all object points. (Similarly, the official instance segmentation evaluation of nuScenes only considers instances with at least 15 points [5].) In order to unify the number of points per cloud in a batch, we pad each point cloud to the max length in the batch.
We further evaluate on objects extracted from KITTI-360 [23]. We use semantic and instance labels from accumulated scans to extract roughly 40k objects in 3 classes as in [54]: car, person, bicycle. We also perform cross-dataset evaluation on SemanticKITTI [2].
For all experiments, we use the conditioning information from objects in the validation set, generating similar views of new objects of the same category.
Object-level metrics.
To evaluate the ability to generate a sensible variant of a real object given only the conditioning information corresponding to the real object, we use two pointset metrics: Chamfer Distance (CD) and Earth Mover’s Distance (EMD).
To evaluate perceptual realism, we use the Fréchet PointNet Distance (FPD) [40], an adaptation of the classic Fréchet Inception Distance (FID) [14], by replacing 2D Inception features with 3D PointNet features [36]. However, the original FPD used as reference relies on the original PointNet classifier, trained on ModelNet40 [36], which is not suited for LiDAR objects. We thus use instead a new PointNet classifier trained on the real LiDAR objects of nuScenes. We additionally consider the Kernel Inception Distance (KID), which has advantages over FID, in particular for small datasets [3]. Like for FID, we adapt KID into a Kernel PointNet Distance (KPD), replacing the Inception network by our trained PointNet.
We also consider distribution comparisons at a pointset level. We use coverage (COV) [1], which measures the proportion of real pointsets matched to at least one generated pointset, and 1-nearest neighbor accuracy (1-NNA) [55], derived from [26], which merges real and synthetic objects and counts the proportion of closest neighbors of each sample that are of the same source. 1-NNA ranges from 100% when distributions are separate, to 50% when they are indistinguishable. COV and 1-NNA are measured with CD and EMD. We extend COV and 1-NNA to measure the quality of the generated intensities. We extract 256-dimensional normalized intensity histograms from generated objects to use as features for COV and 1-NNA computations. Last, in [54], generated objects are extracted from generated scenes after being found by a 3D detector. On KITTI-360, we compare our generations to such detected objects, using object level metrics given in [54]. The Appendix details all metrics.
Model Emb. Point metrics 1-NNA (%)  $\downarrow$ COV (%)  $\uparrow$ FPD $\downarrow$ KPD $\downarrow$
CD EMD CD EMD Int. CD EMD Int. 3 ch. 4 ch. 3 ch. 4 ch.
MinkUNet - 0.208 0.212 79.5 74.4 N/A 33.7 41.7 N/A 25.4 N/A 3.18 N/A
DiT-3DL VE 3 ch 0.136 0.121 75.8 70.9 N/A 34.3 42.0 N/A 2.80 N/A 0.25 N/A
DiT-3DL PE 4 ch 0.142 0.125 73.4 70.7 74.4 35.5 42.1 31.4 2.62 6.99 0.25 0.75
PixArt-L PE 4 ch 0.196 0.204 76.5 72.5 71.6 33.5 41.8 32.9 2.95 5.16 0.24 0.41
LOGen PE 4 ch 0.130 0.111 74.1 70.6 68.2 36.0 42.3 37.0 1.34 2.18 0.10 0.12
Table 1: Combined point-space and feature-space evaluation. CD and EMD are geometric fidelity metrics. 1-NNA and COV (in CD/EMD/intensity space) assess realism and coverage. FPD and KPD evaluate feature distribution alignment under 3- and 4-channel settings (= 3 ch. + intensity). Point embedding (Emb.) is: none (-), voxel-based (VE 3 ch) or PointNet-based (PE 4 ch).
Truck
Refer to caption Refer to caption Refer to caption Refer to caption
Bus
Traffic cone
Refer to caption Refer to caption Refer to caption Refer to caption
Motorbike
Figure 3: Generations with intensities. For each pair, the left object is from the nuScenes val set, the right object is generated using the same conditioning and number of points.
Scene-level metrics.
A second set of evaluations is conducted on the perception of generated objects in a scene context by a typical 3D segmentation model with the contention that more realistic objects should be segmented at a similar rate to real objects. We consider the replacement of ground-truth objects by objects generated with the same conditioning, then measure the change in classwise mIoU segmentation performance of the model on this modified data. We use a variant of nuScenes where each object with at least 20 points is replaced by a generated object of the same class and box size, seen under the same viewing angle, and with the same number of points. We highlight the correspondence between this metric and the #Box metric from [54], where both evaluate the expected performance at identifying generated objects of a model trained to identify real objects.
Baselines.
Due to task novelty, we build baselines: MinkUNet is the sparse convolutional architecture used in LiDiff [34], which ignores intensities. PixArt-L is our adaptation of PixArt-$\alpha$ to LiDAR points and conditioning, using our PointNet embeddings (PE) to allow intensity generation. DiT-3DL VE 3 ch is our adaptation of DiT-3D to LiDAR conditioning, i.e., with its original voxel embeddings (VE) and no intensity handling. (The voxelization library used in DiT-3D [30] does not handle a 4th input channel; handling intensities would require adapting the voxelization layer like in MinkUNet [8], as separate input features.) DiT-3DL PE 4 ch replaces VE by our PE in DiT-3DL VE 3 ch , thus handling intensities.
4.2 Object-level evaluation
Qualitative generation examples are shown in Figs. 1, 3, 4. Fig. 5 presents qualitative comparisons between LOGen and the baseline methods for nuScene classes. In Fig. 6, we present novel objects produced by LOGen.
Point embeddings.
Table 1 shows that the voxel embeddings of DiT-3D do not perform as well as our PointNet embeddings, which supports our proposal (Sect. 3.2). MinkUNet succeeds in generating local structure (LiDAR pattern), but fails to generate objects meeting size conditions. The lack of point embeddings in MinkUNet creates global size inconsistencies.
Architecture.
Table 1 also shows that LOGen transformer blocks perform slightly better than the blocks proposed in PixArt-$\alpha$ and DiT-3D. The superiority of LOGen blocks is more pronounced as measured by the feature-based metrics of FPD and KPD.
Method CD $\downarrow$ JSD $\downarrow$
LiDARGen [60] 33.10 0.41
UltraLiDAR [53] 18.30 0.70
R2DM [32] 17.60 0.86
OLiDM [54] 1.88 0.91
LOGen (ours) 0.24 0.19
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 4: Quantitative & qualitative comparison of LiDAR objects on KITTI-360 [23]. (Top) Evaluation on 1000 sampled objects (for LOGen) or detected objects (for others) using Chamfer Distance (CD) and Jensen-Shannon Divergence (JSD). (Bottom) Examples of generated objects from KITTI-360 for person and car. Classwise results of LOGen and additional metrics EMD, FPD, and KPD are in the appendix.
Comparison on KITTI-360.
The table in Fig. 4 shows the superior visual fidelity of objects generated by LOGen when compared to other scene and object generative methods.
4.3 Scene-level evaluation
Model Emb. barrier bicycle bus car con.veh. moto. pedes. tra.con. trailer truck mean
3 channel MinkUNet - 76.1 31.4 18.7 84.4 14.7 63.6 69.8 59.3 12.4 49.7 48.0
DiT-3DL VE 3 ch 48.4 42.5 0.3 34.1 0.2 63.8 75.6 69.6 7.7 3.4 34.6
DiT-3DL PE 4 ch 61.6 39.2 11.1 84.2 1.7 68.8 77.5 51.3 15.9 26.6 43.9
PixArt-L PE 4 ch 58.9 32.1 3.9 75.7 0.9 76.6 79.0 56.7 13.6 51.3 44.9
LOGen PE 4 ch 78.2 50.9 60.1 79.1 28.1 82.2 84.2 73.5 37.5 63.2 63.7
GT 80.3 48.8 90.7 94.2 40.6 86.0 85.5 73.4 60.7 85.2 74.5
PixArt-L PE 4 ch 56.6 33.7 14.5 82.4 8.1 78.5 79.3 62.8 26.7 52.6 49.5
LOGen PE 4 ch 78.0 48.9 71.7 91.3 37.1 82.0 84.8 79.1 45.1 72.9 69.1
GT 81.6 50.4 90.9 94.6 44.5 88.1 85.6 77.4 64.6 85.6 76.3
Table 2: Scene-level evaluation. SPVCNN is trained with 3 or 4 channels (coordinates + intensity) on nuScenes (train set) with real objects. It is tested on nuScenes (val set) where each object (with $\geq$ 20 points) is replaced with a generated object of the same class, box size and sensor viewing angle, and with the same number of points, and where objects with less than 20 points are removed from the dataset. Displayed results are class-wise IoU%.
Point metrics 1-NNA $\downarrow$ COV $\uparrow$ FPD $\downarrow$ KPD  $\downarrow$ APC  $\uparrow$
LOGen PE 4 ch CD $\downarrow$ EMD $\downarrow$ CD EMD CD EMD 3 ch. 4 ch. 3 ch. 4 ch. 3 ch. 4 ch.
Table 3: Single-model-per-class vs. multi-class model. Comparison of a same size single-model-per-class and multi-class models on object generation metrics. The multi-class model lags behind the single-model-per-class approach in realism metrics.
Semantic quality of generated data.
Table 2 shows that the LOGen generations are more semantically identifiable from the point of view of an SPVCNN segmenter. Though trained on real objects, the segmenter is able to recognize the semantics of LOGen objects at a much higher mIoU than other models, +19.6 pts vs the second best model. This is particularly pronounced for rare classes such as bicycle, bus, trailer, and construction vehicle. Besides, intensities generated by LOGen are realistic enough to provide +5.4 mIoU% at scene level.